# Clinical Data Goes MEDS? Let’s OWL make sense of it

### v1

@@ Abstract
The application of machine learning on healthcare data is often hindered by the lack of standardized and semantically explicit representation, leading to limited interoperability and reproducibility across datasets and experiments. The Medical Event Data Standard (MEDS) addresses these issues by introducing a minimal, event-centirc data model designed for reproducible machine-learning workflows from health data. However, MEDS is defined as a data-format specification and does not natively provide integration with the Semantic Web ecosystem. In this article, we introduce MEDS-OWL, a lightweight OWL ontology that provides formal concepts and relations to enable representing MEDS datasets as RDF graphs. Additionally, we implemented meds2rdf, a Python conversion library that transforms MEDS events into RDF graphs, ensuring conformance with the ontology. We demonstrate the approach on a synthetic clinical dataset that describes patient care pathways for ruptured intracranial aneurysms and validate the resulting graph using a SHACL constraints. The first release of MEDS-OWL comprises 13 classes, 10 object properties, 20 data properties, and 24 OWL axioms. Combined with meds2rdf, it enables data transformation into FAIR-aligned datasets, provenance-aware publishing and interoperability of event-based clinical data. By bridging MEDS with the Semantic Web, this work contributes a reusable semantic layer for event-based clinical data and establishes a robust foundation for subsequent graph-based analytics.
keywords: MEDS, Semantic Web, Knowledge graph, Healthcare interoperability, OWL ontology, FAIR data
@@ 1 Introduction
Machine learning (ML) in the healthcare domain is hindered by the lack of a standardized representation for structured Electronic Health Record (EHR) data, resulting in dataset-specific preprocessing practices, limited reproducibility, and poor model portability across datasets. The Medical Event Data Standard (MEDS) [1, 12] addresses this gap by introducing a unified, event-centric schema for longitudinal health data that is explicitly designed to support reproducible and efficient ML workflows. MEDS encodes each event as a minimal tuple consisting of a subject identifier, a code, and an optional temporal and value attributes. This minimal and pragmatically oriented design makes MEDS particularly suitable for benchmarking, multi-center studies and the development of foundational models. For these reasons, MEDS is beginning to be adopted in various ML projects as a common schema for EHR data and has been utilized in transformer-based model for predicting future events in patient healthcare trajectories [17, 16].
@@
In its native form, MEDS is a data-format specification and does not provide ontology-level semantics. Whereas the Semantic Web paradigm offers a complementary solution [3]. First, ontologies expressed in Resource Description Framework (RDF) and Web Ontology Language (OWL) provide an interoperable, machine-interpretable lingua franca for representing entities and their relationships. Second, they support semantic querying and formal reasoning. Third, they support the construction of knowledge graphs (KGs) that facilitate explicit linkage of dataset elements to external terminologies such as LOINC, SNOMED CT, ATC [4]. Projecting MEDS datasets into RDF would thus enable their semantic integration and validation, while also creating opportunities to explore graph-based ML on event-oriented datasets. Indeed, patient representations can be learned from graph representations with several advantages. KG embeddings map graph nodes and edges into a continuous vector space [2], allowing complex relational structures to be represented in a form amenable to computational analysis. Such approaches have been successfully applied to tasks including link prediction, node classification, and clustering [19]. In the clinical domain, relational models such as Relational Graph Convolutional Networks (RGCNs) have been shown to outperform tabular baselines when the underlying schema aligns with task-specific requirements; for instance, compact and patient-centred representations for outcome prediction [5]. These observations motivate the present work.
@@
While several data models and ontologies, such as FHIR [8], OMOP CDM [21], and Phenopackets [6], have been proposed to harmonize clinical data, RDF transformations of these models have seen limited adoption [21, 9], in part due to loosely specified semantics or incomplete coverage of the clinical domain. Ontologies such as SPHN [18] and CARE-SM [7] illustrate that patient-centric KG representations can facilitate interoperability, provenance modeling, and temporal reasoning. However, these models are often tightly coupled to specific use case or semantic frameworks.
@@
In this study, we propose to bridge MEDS with the Semantic Web formalism. We report on the development of two complementary components: MEDS-OWL, an OWL ontology that captures the core concepts and relations of MEDS [10], and meds2rdf, a Python library that transforms MEDS-compliant datasets into RDF graphs in accordance with the ontology [11]. Together, these components enable interoperable, reproducible and FAIR-aligned representations of event-based clinical data.
@@
The remainder of this article is organized as follows. Section 2 describes the methodology adopted for the development of the MEDS-OWL ontology and the accompanying conversion library. Section 3 presents the resulting ontology and the meds2rdf library. Finally, Section 4 discusses the results and outlines directions for future work.
@@ 2 Methodology
The MEDS-OWL ontology was developed following the principled and iterative methodology described by Noy and McGuinness in “Ontology Development 101” [15], using the Protégé ontology editor [14]. The scope of the ontology encompasses the representation of EHR and claims-style medical events as defined by the MEDS conceptual model, together with the dataset- and experiment-level metadata required to support reproducible ML workflows. As part of this process, we formulated a set of competency questions to guide the modeling choices and to ensure that the ontology adequately supports the intended use cases. An example of such a competency question is: “Given a medical event, which subject does it pertain to, when did it occur, and which canonical code and value modalities are associated with it?”. These competency questions informed the identification and formalization of the core MEDS classes, namely Subject, Event, Code, DatasetMetadata, SubjectSplit, and SubjectLabel. A detailed description of these six core classes, together with illustrative examples, is provided in Table 1.
@@
Code Entity used to type events, which includes canonical identifiers, descriptions, and hierarchical relations, with optional linkage to external terminologies. ATC:C08CA06 (Nimodipine)
@@
SubjectSplit A grouping of subjects in several disjoint sets aimed for a ML experiment. Split for experiment 5 in a train, eval and test sets
@@
The OWL classes and properties we developed follow closely the MEDS data model, while reusing as much as possible established external ontologies to maximize interoperability. Specifically, we reused the Data Catalogue Vocabulary (DCAT), DCTERMS, and PROV ontology (PROV-O) to represent datasets, ETL (Extraction, Transformation, Loading) processes, and provenance metadata, respectively, thereby ensuring alignment with widely adopted ontologies. Finally, for MEDS fields specified as optional (nullable), such as time, the ontology captures this constraint through upper-bound cardinality restrictions (owl:maxCardinality 1) without corresponding minimum bound or existential constraints, thus permitting zero or one value, while forbidding multiple values.
@@
We developed a Python library named meds2rdf [11] to implement the mapping rules necessary to transform a MEDS dataset into an RDF graph that follows the terms of the MEDS-OWL ontology. Figure 1 presents the end-to-end conversion workflow implemented leveraging meds2rdf. The pipeline consists of three main stages. (a) First, source EHR data are ingested by the MEDS-extract component [12], which applies a declarative conversion template to map dataset-specific features into MEDS event tuples. Note that this first step is optional, and only necessary when healthcare data are not already available in MEDS. (b) In the subsequent RDF generation stage, MEDS-formatted data are transformed into a KG conforming to the MEDS-OWL ontology using the meds2rdf library. The mapping module of meds2rdf implements a deterministic, table-driven transformation from MEDS tuples to RDF. It exploits manually curated mapping tables that relate elements of the MEDS schema to corresponding classes and properties in MEDS-OWL. These tables constitute a central contribution of this work, as they provide an implementation-independent and ontology-aligned specification that formalizes the MEDS to RDF mapping. Moreover, in this phase, meds2rdf focuses on syntactic correctness and local mapping guarantees, such as required-field checks and type normalization. The output of this stage is an in-memory RDF graph constructed using the rdflib.Graph data structure. (c) Next, the generated graph is validated against a SHACL schema using the pySHACL validator, which is integrated within meds2rdf. The schema declaratively encodes the semantic constraints of MEDS-OWL, and the validation process ensures the enforcement of both global and cross-entity invariants. These include mutually exclusive property patterns, cardinality constraints, and referential consistency across entities. This approach allows semantic validation to be performed prior to graph materialization. Finally, RDF graphs that successfully pass SHACL validation may be serialized as an optional final step in standard RDF formats, including Turtle, RDF/XML, and N-Triples.
Integrating SHACL validation within meds2rdf, rather than implementing these constraints programmatically in the conversion code, enables the ontology and its constraints to evolve independently of the mapping logic. This design choice is particularly important given that MEDS-OWL is currently at its first public version and is expected to undergo further revisions in alignment with the evolving MEDS specification. By leveraging the declarative nature of SHACL, future updates to the ontology or constraint definitions can be accommodated by modifying the SHACL schema alone, without requiring changes to the conversion code.
3 Results and example usage
@@
Mappings between MEDS core data schema elements and MEDS-OWL are reported in Tables 2-5. Table 2 reports those to the meds:Event class. In MEDS-OWL, events are required to be associated with exactly one subject and exactly one code, while temporal and value-related attributes are optional and for this reason constrained by cardinality restrictions. Table 3 reports mappings to the meds:Code class. Code metadata is modeled explicitly as individuals of type meds:Code, which enables reusability, hierarchical organization, and linkage to external terminologies. Dataset-level provenance and conversion metadata are represented by the meds:DatasetMetadata class, which is aligned with both dcat:Dataset and prov:Entity. Cardinality constraints are enforced at the OWL level. The corresponding mappings are shown in Table 4. Finally, Table 5 reports the mapping for dataset splits and supervised learning labels. Dataset partitions are modeled as predefined individuals of class meds:SubjectSplit, while label samples are represented as instances of meds:SubjectLabel with exactly one label value. The meds2rdf library that implements the mapping rules is available on Zenodo and gitHub [11].
@@
subject_id id The identifer of the subject s (typically the patient). ?s a meds:Subject. ?s meds:subjectId ?id^^xsd:string. ?e meds:hasSubject ?s. 1
@@
Table 4: Mapping metadata of a transformed MEDS dataset to the meds:DatasetMetdata class. For each dataset generated, a list of associated “meta” RDF triples is generated.
@@
To example the usage of meds2rdf, we employed a synthetic EHR dataset, here named neurovasc, that describes care pathways of 10,000 patients hospitalized for a ruptured intracranial aneurysm [5]. neurovasc is an open dataset generated based on real-world EHR from Nantes University Hospital. Its patients are associated with 30 clinical features, including 8 time-stamped events representing the interval from hospital admission to their first occurrence, and 22 non-temporal demographic or historical features. Patient outcomes take one of three values: Back home, Rehabilitation, or Death, depending on the status of the patient at the end of the hosptial stay. The dataset preserves realistic correlations and event transitions, making it suitable for testing ontology-driven data representation of clinical pathways and basics analytics.
The conversion of the neurovasc dataset illustrates the practical utility of MEDS-OWL. As shown in Figure 3, the workflow depicts the output of each conversion stage, starting from a simplified representation of the original neurovasc data. To provide visual clarity, only three patient features are shown: age, administration of nimodipine (ATC code: C08CA06), and a tracheostomy procedure (ICD-10-PCS code: 0BH17EZ). (a) Following an initial preprocessing step, (b) the data are passed to MEDS-extract and converted into the MEDS tabular format, where each feature is represented as a distinct event optionally linked to external code systems. (c) In the final stage, the MEDS root folder containing the ETL output is provided as input to the meds2rdf library, which systematically transforms MEDS events into RDF graphs by instantiating the corresponding MEDS-OWL ontology classes. The resulting graphs preserve source information and enable semantic validation through SHACL shapes. The exemplified used of our tools to transform neurovasc to MEDS and then to RDF are publicly available at https://github.com/albertomarfoglia/neurovasc_on_meds.
@@
To provide a more objective characterization of the transformation process and the resulting KG, we report quantitative statistics at each stage of the neurovasc pipeline shown in Figure 3. The original neurovasc dataset contains 10,000 rows, each describing 30 patient-level features. After preprocessing and conversion with MEDS-extract, the dataset is transformed into 245,015 MEDS events, reflecting the expansion of patient-level features into event-based representations. These events are distributed across the standard ML splits as follows: 196,064 events in the train set, 24,474 in the validation set, and 24,477 in the test set.
@@
Finally, we analyzed the structural footprint of events in the resulting RDF graph. Each meds:Event node is described, on average, by 5.31 RDF triples (standard deviation 0.46), with a minimum of 5 and a maximum of 6 triples per event. This low variance reflects the regular structure imposed by MEDS-OWL cardinality constraints, where subject and code associations are mandatory, while temporal and value-related properties are optional. These statistics demonstrate that the transformation produces a compact RDF representation of events, while preserving semantic richness and provenance information.

### v2

@@ Abstract
The application of machine learning on healthcare data is often hindered by the lack of standardized and semantically explicit representation, leading to limited interoperability and reproducibility across datasets and experiments. The Medical Event Data Standard (MEDS) addresses these issues by introducing a minimal, event-centric data model designed for reproducible machine-learning workflows from health data. However, MEDS is defined as a data-format specification and does not natively provide integration with the Semantic Web ecosystem. In this article, we introduce MEDS-OWL, a lightweight OWL ontology that provides formal concepts and relations to represent MEDS datasets as RDF graphs. Additionally, we implemented meds2rdf, a Python conversion library that transforms MEDS events into RDF graphs, ensuring conformance with the ontology. We evaluate the proposed approach on two datasets: a synthetic clinical cohort describing care pathways for ruptured intracranial aneurysms, and a real-world subset of MIMIC-IV. To assess semantic consistency, we performed a SHACL validation against the resulting knowledge graphs. The first release of MEDS-OWL comprises 13 classes, 10 object properties, 20 data properties, and 24 OWL axioms. Combined with meds2rdf, it enables data transformation into FAIR-aligned datasets, provenance-aware publishing, and interoperability of event-based clinical data. By bridging MEDS with the Semantic Web, this work contributes a reusable semantic layer for event-based clinical data and establishes a robust foundation for subsequent graph-based analytics.
@@ 1 Introduction
Machine learning (ML) in the healthcare domain is hindered by the lack of a standardized representation for structured Electronic Health Record (EHR) data, resulting in dataset-specific preprocessing practices, limited reproducibility, and poor model portability across datasets. The Medical Event Data Standard (MEDS) [1, 16] addresses this gap by introducing a unified, event-centric schema for longitudinal health data that is explicitly designed to support reproducible and efficient ML workflows. MEDS encodes each event as a minimal tuple consisting of a subject identifier, a code, and an optional temporal and value attributes. This minimal and pragmatically oriented design makes MEDS particularly suitable for benchmarking, multi-center studies, and the development of foundational models. For these reasons, MEDS is being adopted in various ML projects as a common schema for EHR data and has been utilized in a transformer-based model for predicting future events in patient healthcare trajectories [21, 20].
@@
In its native form, MEDS is a data-format specification and does not provide ontology-level semantics. The Semantic Web paradigm offers a complementary approach to address this limitation [4, 2, 12]. First, ontologies expressed in Resource Description Framework (RDF) and Web Ontology Language (OWL) provide an interoperable, machine-interpretable lingua franca for representing entities and their relationships. Second, they support semantic querying and formal reasoning. Third, they support the construction of knowledge graphs (KGs) that facilitate the explicit linkage of dataset elements to external terminologies such as LOINC, SNOMED CT, and ATC [5]. Projecting MEDS datasets into RDF would thus enable their semantic integration and validation, while also creating opportunities to explore graph-based ML on event-oriented datasets. Indeed, patient representations can be learned from graph representations with several advantages. KG embeddings map graph nodes and edges into a continuous vector space [3], enabling the representation of complex relational structures in a form amenable to computational analysis. Such approaches have been successfully applied to tasks including link prediction, node classification, and clustering [23]. In the clinical domain, relational models such as Relational Graph Convolutional Networks (RGCNs) have been shown to outperform tabular baselines when the underlying schema aligns with task-specific requirements; for instance, compact and patient-centered representations for outcome prediction [6]. These observations motivate the present work.
@@
While several data models and ontologies, such as FHIR [10], OMOP CDM [24], and Phenopackets [8], have been proposed to harmonize clinical data, RDF transformations of these models have seen limited adoption [24, 13], in part due to loosely specified semantics or incomplete coverage of the clinical domain. Patient-centric RDF-based representations such as SPHN [22] and CARE-SM [9] illustrate how KGs can facilitate interoperability, provenance modeling, and temporal reasoning. However, these models are often tightly coupled to specific use cases or semantic frameworks.
@@
In this study, we propose to bridge MEDS with the Semantic Web formalism. We report on the development of two complementary components: MEDS-OWL, an OWL ontology that captures the core concepts and relations of MEDS [14], and meds2rdf, a Python library that transforms MEDS-compliant datasets into RDF graphs in accordance with the ontology [15]. Together, these components enable interoperable, reproducible, and FAIR-aligned representations of event-based clinical data.
@@
The remainder of this article is organized as follows. Section 2 describes the methodology adopted for the development of the MEDS-OWL ontology and the accompanying conversion library. Section 3 presents the resulting ontology and the meds2rdf library in action. Finally, Section 4 discusses the results and outlines directions for future work.
@@ 2 Methodology
The MEDS-OWL ontology was developed following the principled and iterative methodology described by Noy and McGuinness in “Ontology Development 101” [19], using the Protégé ontology editor [18]. The scope of the ontology encompasses the representation of EHR and claims-style medical events as defined by the MEDS conceptual model, together with the dataset- and experiment-level metadata required to support reproducible ML workflows. As part of this process, we formulated a set of competency questions to guide the modeling choices and to ensure that the ontology adequately supports the intended use cases. An example of such a competency question is: “Given a medical event, which subject does it pertain to, when did it occur, and which canonical code and value modalities are associated with it?”. These competency questions informed the identification and formalization of the core MEDS classes, namely Subject, Event, Code, DatasetMetadata, SubjectSplit, SubjectLabel, and ValueModality. A detailed description of these Seven core classes, together with illustrative examples, are provided in Table 1.
@@
Code An entity used to type events, which includes canonical identifiers, descriptions, and hierarchical relations, with optional linkage to external terminologies. ATC:C08CA06 (Nimodipine)
@@
SubjectSplit A grouping of subjects in several disjoint sets aimed for an ML experiment. Split for experiment 5 into a train, eval, and test sets
@@
ValueModality Generic, extensible entity that links an Event to additional value modalities (for example, images, waveform recordings, audio clips, or other non-tabular payloads). Chest X-ray image (DICOM series URI)
To clarify how these ontology classes relate to the original MEDS data model, Table 2 summarizes the correspondence between the five primary MEDS schema components and the MEDS-OWL classes derived from them. In MEDS, each schema is defined as an Apache Arrow schema, specifying the required columns and data types for MEDS data files.
Table 2: Mapping from MEDS schema components to MEDS-OWL classes. See Table 1 for class definitions and examples.
MEDS schema Brief description Mapped MEDS-OWL class(es)
DataSchema Core medical event table describing sequences of subject observations, where each row corresponds to a single clinical event. meds:Event as the primary representation of each MEDS row. meds:ValueModality is used to model additional or non-tabular payloads associated with an event (e.g., images, waveforms, signals). See Table 3 for row-to-triple mappings.
DatasetMetadataSchema Dataset-level metadata captures dataset identity, versioning, creation time, licensing, and ETL provenance information. meds:DatasetMetadata, aligned with dcterms and dcat, and linked to ETL activities via prov:wasGeneratedBy. See Table 5.
CodeMetadataSchema Metadata describing the controlled vocabularies and codes used to type medical events, including canonical identifiers and hierarchical relations. meds:Code, with meds:codeString as a canonical identifier and optional meds:parentCode relations to represent code hierarchies. See Table 4.
SubjectSplitSchema Specification of how subjects are assigned to mutually exclusive subsets (e.g., training, tuning, held-out) within a given machine learning experiment. meds:SubjectSplit, with subjects linked via meds:assignedSplit. Multiple splits may coexist across experiments, while each subject belongs to exactly one subset per split. See Table 6.
LabelSchema Definition of supervised learning labels associated with a subject at a specific prediction time, including the observed ground truth value. meds:SubjectLabel, capturing the prediction cutoff time (meds:predictionTime) and exactly one label value using datatype-specific properties. See Table 6.
(subject identifier) Subject identifiers appear across multiple MEDS schemas and provide the logical join key between events, splits, and labels. meds:Subject, representing the canonical subject entity via meds:subjectId, and serving as the hub linking Event, SubjectLabel, and SubjectSplit.
The OWL classes and properties we developed closely follow the MEDS data model, while reusing as much as possible established external ontologies to maximize interoperability. Specifically, we reused the Data Catalogue Vocabulary (DCAT), DCTERMS, and PROV ontology (PROV-O) to represent datasets, ETL (Extraction, Transformation, Loading) processes, and provenance metadata, respectively, thereby ensuring alignment with widely adopted ontologies. Finally, for MEDS fields specified as optional (nullable), such as time, the ontology captures this constraint through upper-bound cardinality restrictions (owl:maxCardinality 1) without corresponding minimum bound or existential constraints, thus permitting zero or one value, while forbidding multiple values.
@@
We developed a Python library named meds2rdf [15] to implement the mapping rules necessary to transform a MEDS dataset into an RDF graph that follows the terms of the MEDS-OWL ontology. Figure 1 presents the end-to-end conversion workflow implemented leveraging meds2rdf. The pipeline consists of three main stages. (a) First, source EHR data are ingested by the MEDS-extract component [16], which applies a declarative conversion template to map dataset-specific features into MEDS event tuples. Note that this first step is optional, and only necessary when healthcare data are not already available in MEDS. (b) In the subsequent RDF generation stage, MEDS-formatted data are transformed into a KG conforming to the MEDS-OWL ontology using the meds2rdf library. The mapping module of meds2rdf implements a deterministic, table-driven transformation from MEDS tuples to RDF. It exploits manually curated mapping tables that relate elements of the MEDS schema to corresponding classes and properties in MEDS-OWL. These tables constitute a central contribution of this work, as they provide an implementation-independent and ontology-aligned specification that formalizes the MEDS-to-RDF mapping. Moreover, in this phase, meds2rdf focuses on syntactic correctness and local mapping guarantees, such as required-field checks and type normalization. The output of this stage is an in-memory RDF graph constructed using the rdflib.Graph data structure. (c) Next, the generated graph is validated against a SHACL schema using the pySHACL validator, which is integrated within meds2rdf. The schema declaratively encodes the semantic constraints of MEDS-OWL, and the validation process ensures the enforcement of both global and cross-entity invariants. These include mutually exclusive property patterns, cardinality constraints, and referential consistency across entities. This approach allows semantic validation to be performed before graph materialization. Finally, RDF graphs that successfully pass SHACL validation may be serialized as an optional final step in standard RDF formats, including Turtle, RDF/XML, and N-Triples.
Integrating SHACL validation within meds2rdf enables the ontology and its constraints to evolve independently of the mapping logic. This design choice is crucial given that MEDS-OWL is currently at its first public version and is expected to undergo further revisions in alignment with the evolving MEDS specification. By leveraging the declarative nature of SHACL, future updates to the ontology or constraint definitions can be accommodated by modifying the SHACL schema alone, without requiring changes to the conversion code.
3 Results and examples of usage
@@
Mappings between MEDS core data schema elements and MEDS-OWL are reported in Tables 3-6. Table 3 reports those to the meds:Event class. In MEDS-OWL, events are required to be associated with exactly one subject and exactly one code, while temporal and value-related attributes are optional and for this reason constrained by cardinality restrictions. Table 4 reports mappings to the meds:Code class. Code metadata is modeled explicitly as individuals of type meds:Code, which enables reusability, hierarchical organization, and linkage to external terminologies. Dataset-level provenance and conversion metadata are represented by the meds:DatasetMetadata class, which is aligned with both dcat:Dataset and prov:Entity. Cardinality constraints are enforced at the OWL level. The corresponding mappings are shown in Table 5. Finally, Table 6 reports the mapping for dataset splits and supervised learning labels. Dataset partitions are modeled as predefined individuals of class meds:SubjectSplit, while label samples are represented as instances of meds:SubjectLabel with exactly one label value. The meds2rdf library that implements the mapping rules is available on Zenodo and GitHub [15].
@@
subject_id id The identifier of the subject s (typically the patient). ?s a meds:Subject. ?s meds:subjectId ?id^^xsd:string. ?e meds:hasSubject ?s. 1
@@
Table 5: Mapping metadata of a transformed MEDS dataset (DatasetMetadataSchema) to the meds:DatasetMetdata class. For each dataset generated, a list of associated “meta” RDF triples is generated.
@@
Neurovasc example of usage. To evaluate meds2rdf, we first employed a synthetic EHR dataset, here named neurovasc, that describes care pathways of 10,000 patients hospitalized for a ruptured intracranial aneurysm [6]. neurovasc is an open dataset generated based on real-world EHR from Nantes University Hospital. Its patients are associated with 30 clinical features, including 8 time-stamped events representing the interval from hospital admission to their first occurrence, and 22 non-temporal demographic or historical features. Patient outcomes are categorized in one of three values: Back Home, Rehabilitation, or Death, depending on the patient’s status at the end of the hospital stay. The dataset preserves realistic correlations and event transitions, making it suitable for testing ontology-driven data representation of clinical pathways and basic analytics.
The conversion of the neurovasc dataset illustrates the practical utility of MEDS-OWL. As shown in Figure 3, the workflow depicts the output of each conversion stage, beginning with a simplified representation of the original neurovasc data. To provide visual clarity, only three patient features are shown: age, administration of nimodipine (ATC code: C08CA06), and a tracheostomy procedure (ICD-10-PCS code: 0BH17EZ). (a) Following an initial preprocessing step, (b) the data are passed to MEDS-extract and converted into the MEDS tabular format, where each feature is represented as a distinct event optionally linked to external code systems. (c) In the final stage, the MEDS root folder containing the ETL output is provided as input to the meds2rdf library, which systematically transforms MEDS events into RDF graphs by instantiating the corresponding MEDS-OWL ontology classes. The resulting graphs preserve source information and enable semantic validation through SHACL shapes. The exemplified use of our tools to transform neurovasc to MEDS and then to RDF is publicly available at https://github.com/albertomarfoglia/meds-to-owl-examples.
@@
To provide a more objective characterization of the transformation process and the resulting KG, we report quantitative statistics at each stage of the neurovasc pipeline shown in Figure 3. The original dataset contains 10,000 rows, each describing 30 patient-level features. After preprocessing and conversion with MEDS-extract, the dataset is transformed into 245,015 MEDS events, reflecting the expansion of patient-level features into event-based representations. These events are distributed across the standard ML splits as follows: 196,064 events in the train set, 24,474 in the validation set, and 24,477 in the test set.
@@
Finally, we analyzed the structural footprint of events in the resulting RDF graph. Each meds:Event node is described, on average, by 5.31 RDF triples (standard deviation 0.46), with a minimum of 5 and a maximum of 6 triples per event. This low variance reflects the regular structure imposed by MEDS-OWL cardinality constraints, where subject and code associations are mandatory, while temporal and value-related properties are optional. These statistics demonstrate that the transformation produces a compact RDF representation of events while preserving semantic richness and provenance information.
MIMIC-IV example of usage. To further demonstrate the scalability and applicability of meds2rdf to real-world EHRs, we applied the library to a subset of the publicly available MIMIC-IV demo dataset [7] already formatted in MEDS [17]. MIMIC-IV is a large, openly accessible critical care dataset containing de-identified EHRs for ICU patients from Beth Israel Deaconess Medical Center. It is widely used in clinical informatics research for benchmarking predictive models, phenotyping, and analyzing temporal data. For this experiment, we selected a subset of 100 patients to illustrate the transformation process while keeping the workflow tractable for visualization and reproducibility.
The conversion generated 916,166 MEDS events distributed across the ML splits as follows: 803,992 in the train set, 69,620 in the tuning set, and 42,554 in the held-out set. The resulting KG comprises 6,408,529 RDF triples, including 925,706 distinct subjects, 19 predicates, 126,680 distinct objects, 925,733 distinct IRIs, and 117,132 literals, with no blank nodes generated. Ontology instantiation includes 916,166 individuals of class meds:Event, 100 meds:Subject individuals, 9,435 meds:Code individuals, three meds:SubjectSplit individuals, and one meds:DatasetMetadata individual. No label or ValueModality instances were included in this subset.
Each meds:Event in the MIMIC-IV subset is represented by an average of 6.97 RDF triples (median 7.0, minimum 5, maximum 8, standard deviation 0.42), indicating a slightly richer structure per event compared to neurovasc, due to the higher number of numeric and text values associated with MIMIC events. These results confirm that meds2rdf can handle real-world MEDS datasets efficiently while preserving the semantic and provenance structure defined by MEDS-OWL.